Find the most general antiderivative for the function 6x−−√3−5x2−−√3.
Antiderivative =
1
Expert's answer
2020-12-15T01:19:02-0500

"\begin{aligned}
\int 6x− √3 − 5x^2 − √3\, \mathrm{d}x &= 3x^2 - \sqrt{3}x - \sqrt{3}x - \frac{5x^3}{3} + C
\\&= 3x^2 - 2\sqrt{3}x- \frac{5x^3}{3} + C
\end{aligned}"